# Emergent vorticity asymmetry of one and two-layer shallow water system captured by a next-order balanced model

### v3

@@ 2.2.1 Set-ups of the simulations
The initial conditions are carefully constructed to benefit the study of vortex asymmetry. They should have zero vorticity skewness initially and be well-balanced to avoid excessive gravity-wave emission. For this, we use the procedure based on the Balance Equations used in Polvani et al. (1994, (2.5)). The vorticity is set to be isotropic Gaussian random fields (thus no skewness) centered around nondimensional size $4$ and nondimensional wavenumber $1.6$, crudely mimicking energy injected by baroclinic instability. The magnitude is determined so that the QG eddy kinetic energy of the vorticity field is one-half:
@@
The fixed-point algorithm described in Polvani et al. (1994) is then performed to solve for the well-balanced divergence and height fields to match the vorticity fields.
Vorticity asymmetry emerges from the evolution of the shallow water model. We evolve the simulation pseudospectrally using Dedalus (Burns et al., 2020) for 500 nondimensional time. The horizontal resolution is $1024\times 1024$ Fourier modes. We dissipate the small-scale for numerical stability using a fourth-order hyper-dissipation just large enough to absorb small-scale noise, as diagnosed by the vorticity spectrum
$\nu\nabla^{4}(\cdot)\quad\mbox{with {\color[rgb]{0,0,0}non-dimensional value}}\quad\nu=5.5\times 10^{-6}.$ (32)
on $u,v,h$ for the shallow water simulation and on $q$ for the SWQG+1 simulation. Time-stepping uses the 3rd-order 4-stage diagonally-implicit+explicit Runge-Kutta (DIRK+ERK) scheme (Ascher et al., 1997). The timestep size is determined to have a CFL number of 0.5.
@@ 2.2.2 Simulation results
Ensemble statistics are appropriate for studying the emergent vorticity asymmetry of the shallow water model and whether the SWQG+1 model can capture it. The left of Figure 2 shows the time evolution of vorticity skewness of the size ten ensemble of simulations of the shallow water model and SWQG+1, for $\varepsilon=0.1$. The initial conditions are constructed to have symmetric vorticity, and indeed, the vorticity skewness starts at zero for all simulations. All simulations’ vorticity skewness trends towards negative. The spread between members of the ensembles becomes larger over time, as expected for the chaotic dynamics of shallow water turbulence. However, the ensemble means of vorticity skewness of the two models agree remarkably well during the entire time series. They lie within $1/\sqrt{10}$ times the ensemble standard deviation from each other for most of the time series. We use this range since it is the standard scaling of error of a Monte-Carlo estimate of the ensemble mean. The close agreement indicates strongly that SWQG+1 model can capture the mechanism that leads to vorticity asymmetry in the shallow water model (Arai & Yamagata, 1994; Graves et al., 2006).
@@
The agreement extends to all Rossby numbers explored. The right of Figure 2 shows the ensemble mean and $1/\sqrt{10}$ times the ensemble standard deviation of vorticity skewness for the shallow water model and SWQG+1 at nondimensional time $t/T=200$. We see that vorticity skewness becomes more negative as the Rossby number gets larger. Polvani et al. (1994, FIG. 10) shows a similar trend except with the Froude number, which is equal to the Rossby number for all our simulations where the Burger number is one.
@@
which is equal to the Froude number squared. This is consistent with the fact that vorticity skewness is proportional to the Froude number (Polvani et al., 1994).
@@
The freely evolving simulation in the shallow water model is a simple setting where we can investigate the properties of the SWQG+1 model. We have not been able to show that the SWQG+1 model conserves energy. Instead, we diagnose the energy and show that it mimics the behavior of the energy of the shallow water model. The left of Figure 4 shows the time series of the total energy as well as the EKE and APE components (9) for the evolution of the $\varepsilon=0.1$ ensemble. The EKE decreases as it is converted to the APE, a signature of the inverse cascade. The total energy is well-conserved such that more than 90% of it is in the final state. The SWQG+1 total energy evolution closely follows the one for the full shallow water model, and this is true for all Rossby numbers explored (not shown). However, it is noticeable that the SWQG+1 loses more energy, compared to the shallow water model. The shallow water model dissipates the physical variables separately, while the SWQG+1 model dissipates the PV. Dissipation in the shallow water model does not translate to dissipation of the PV, because of the nonlinearity in the shallow water PV (6). This is well-known for the Ertel PV of the Boussinesq system, where the dissipation of the physical variables translates to a source term in the PV equations (Haynes & McIntyre, 1987). We will not explore the possibility of including source terms in the PV equation in this work any further. The total energy is minimally dissipated, and the dissipation effects are not significant.
@@ 3.1 The two-layer shallow water system and the derivation of the two-layer SWQG+1 model
+$\qquad+\{\varepsilon\}^{2}\left[\left\{\textit{Bu}^{-1}\right\}^{2}\frac{fh_{i}^{2}}{H^{2}}-\left\{\textit{Bu}^{-1}\right\}\frac{(\zeta_{i}+\left\{\gamma\right\}\beta y)h_{i}}{H}\right]+O\left(\{\varepsilon\}^{3}\right)$ (40c)
@@
$\mathcal{L}\begin{pmatrix}\Phi_{1}^{0}\vspace{3mm}\\ \Phi_{2}^{0}\end{pmatrix}=\begin{pmatrix}[l]\nabla^{2}\Phi_{1}^{0}+\left\{\textit{Bu}^{-1}\right\}\frac{f^{2}}{g^{\prime}H}(\Phi^{0}_{2}-\Phi^{0}_{1})-\left\{\textit{Bu}^{-1}\right\}\frac{f^{2}}{gH}\Phi_{1}^{0}\vspace{1mm}\\ \nabla^{2}\Phi_{2}^{0}+\left\{\textit{Bu}^{-1}\right\}\frac{f^{2}}{g^{\prime}H}(\Phi^{0}_{1}-\Phi^{0}_{2})\end{pmatrix}=\begin{pmatrix}[l]q_{1}\vspace{3mm}\\ q_{2}\end{pmatrix}-\left\{\gamma\right\}\beta y.$ (42)
@@
$\mathcal{L}\begin{pmatrix}\Phi_{1}^{1}\vspace{3mm}\\ \Phi_{2}^{1}\end{pmatrix}=\begin{pmatrix}[l]C_{q,1}-\left[\left\{\textit{Bu}^{-1}\right\}^{2}\frac{f}{H^{2}}\left(\frac{f}{g^{\prime}}(\Phi^{0}_{1}-\Phi^{0}_{2})+\frac{f}{g}\Phi^{0}_{1}\right)^{2}\right.\\ \left.\qquad-\left\{\textit{Bu}^{-1}\right\}\frac{1}{H}(\nabla^{2}\Phi^{0}_{1}+\left\{\gamma\right\}\beta y)\left(\frac{f}{g^{\prime}}(\Phi^{0}_{1}-\Phi^{0}_{2})+\frac{f}{g}\Phi^{0}_{1}\right)\right]\vspace{1mm}\\ C_{q,2}-\left[\left\{\textit{Bu}^{-1}\right\}^{2}\frac{f}{H^{2}}\left(\frac{f}{g^{\prime}}(\Phi^{0}_{2}-\Phi^{0}_{1})\right)^{2}\right.\\ \left.\qquad-\left\{\textit{Bu}^{-1}\right\}\frac{1}{H}(\nabla^{2}\Phi^{0}_{2}+\left\{\gamma\right\}\beta y)\left(\frac{f}{g^{\prime}}(\Phi^{0}_{2}-\Phi^{0}_{1})\right)\right]\end{pmatrix}$ (43)
@@
[…]
@@
[…]
@@ 3.2.1 Set-ups of the simulations
[…]
@@ 3.2.2 Simulation results
The SWQG+1 captures the vorticity evolution of the baroclinic life cycle to remarkable accuracy. Figure 7 displays snapshots of the barotropic ($(\zeta_{1}+\zeta_{2})/2$) and baroclinic ($(\zeta_{1}-\zeta_{2})/2$) vorticity (in color) and divergence (in contour) during the growth of the baroclinic instability. Note that the time of the snapshot is offset. This is because the growth of the instability from small-scale noise happens at different times in the model, due to the chaotic nature of the instability. However, the growth rate is remarkably similar (see Figure 8). The match is indeed impressive. Not only does the SWQG+1 model capture the asymmetric evolution of the cyclonic and anticyclonic part of the baroclinic wave, it is also able to recover the associated divergence fields. Note that one could infer the same divergence, using the $\omega$-equation. The difference of SWQG+1 is that the full velocity participates in the advection, while in SWQG, the higher-order divergence is implicit and is not used in the advection of PV. The divergence is of crucial atmospheric interest as a convergent region in the lower layer corresponds to precipitation. For our example, baroclinic convergence correlates with baroclinic cyclonic vorticity, and baroclinic divergence is large near regions of barotropic strain. On the other hand, barotropic divergence is uniformly small. This is reminiscent of the strain-induced fronts configuration studied in Hoskins & Bretherton (1972). We will explore further strain-induced fronts in Section 3.3. The similarity of the evolution for the two models go beyond this early snapshot. In the supplementary materials, we provide videos of the $\{\varepsilon\}=0.1$ simulations.
@@
Figure 7: The vorticity field of the evolution of the jets for the shallow water model (left) and SWQG+1 model (right). The contour is divergence $\{\varepsilon\}\delta/f$ at $[-0.1,-0.005,0.005,0.1]$, with the solid being positive. The shallow water snapshot is taken at $t/T=30$ while the SWQG+1 snapshot is taken at $t/T=24$.
@@
The match of the vorticity snapshots translates to the ability of the SWQG+1 to capture the vorticity skewness of the simulation. Figure 8 shows the agreement of the overall trend over a large time ($t/T=100$), for both layers. The vorticity is now skewed towards positive, in comparison to the freely decaying result. This change in vorticity skewness is due to the new effects of the baroclinic mode now included in the two-layer model. The vortex stretching due to the correlation of baroclinic vorticity and divergence drives the initial growth of positive vorticity skewness shown in Figure 8. This effect is entirely missing in the one-layer simulation. The vorticity skewness is an ageostrophic phenomenon. We diagnose the first local maxima of the vorticity skewness, as it is an indicator of the strength of the ageostrophic frontogenesis. Figure 9 shows clearly that this measurement of vorticity skewness is linearly proportional to the Rossby number, and SWQG+1 is effective over a large parameter space. However, it is worth noting that this close match between shallow water and SWQG+1 might not hold across all Burger numbers. Our chosen regime has minimal spontaneous gravity wave emission, even for high Rossby number. This might not be the case for another choice of Burger number. We leave further exploration of other parameters to further study.
@@
Our results agree with those of Lambaerts et al. (2012) in the upper layer for the onset of instability. However, their lower layer does not have a mean jet while ours does, thus the results differ.
@@ 3.3 Baroclinic strain-driven fronts modeled by SWQG+1
The two-layer simulation reveals that strain-driven fronts are important in the dynamics and contribute to the positive vorticity skewness, which is opposite to the results of our one-layer simulations. This is a new feature of the two-layer shallow water simulations and worth further investigation. Here we study a minimal model of strain-driven fronts by adapting the set-up of Hoskins & Bretherton (1972) to the two-layer.
@@
$\mathcal{L}\begin{pmatrix}G^{1}_{1}\vspace{3mm}\\ G^{1}_{2}\end{pmatrix}$ =$=\left\{\textit{Bu}^{-1}\right\}\frac{2\alpha f}{g^{\prime}H}\begin{pmatrix}[l](\Phi^{0}_{2}-\Phi^{0}_{1})_{,y}\vspace{2mm}\\ (\Phi^{0}_{1}-\Phi^{0}_{2})_{,y}\end{pmatrix}.$ (50)
@@
The Hoskins & Bretherton (1972) model is posed in an infinite domain in $y$, where perturbation velocities decay at infinity. This implies boundary conditions for the potentials:
@@
$\delta_{1}=v_{1,y}$ =$=-\{\varepsilon\}G_{1,y}\quad\mbox{and}\quad\delta_{2}=v_{2,y}=-\{\varepsilon\}G_{2,y}.$ (53)
overlaid on the height perturbations. In the top layer, a cold filament has cyclonic vorticity, which is correlated with convergence. This correlation is a key feature of fronts and explains the cyclonic bias of vorticity in the two-layer simulations by appealing to vortex stretching: baroclinic cyclonic vorticity is correlated with baroclinic negative divergence. Note that the above equation is equivalent to the “$\omega$-equation”. Therefore, it is implicit in the SWQG model. However, SWQG does not use this ageostrophic velocity in its advective evolution. It does not capture the vorticity skewness evolution.
Finally, a comparison with the surface buoyancy gradient driven Boussinesq frontogenesis of Hoskins & Bretherton (1972) is warranted. In the layered shallow water set-up, the convergence will smooth away the height perturbation and release the stored APE. Shallow water fronts would be less strong with no finite time singularity, compared to the Boussinesq case.
Figure 10: Divergence of a cold filament driven by a barotropic strain, modeled by a two-layer rigid lid SWQG+1 model. The domain is divided vertically into two layers by the height perturbation. The color shows divergence values. This is a zoomed-in view showing $y/L=[-3,3]$

### v4

@@ 2.2.1 Set-ups of the simulations
The initial conditions are carefully constructed to benefit the study of vortex asymmetry. They should have zero vorticity skewness initially and be well-balanced to avoid excessive gravity-wave emission. For this, we use the procedure based on the Balance Equations used in PolvaniEtAl_94. The vorticity is set to be isotropic Gaussian random fields (thus no skewness) centered around nondimensional size $4$ and nondimensional wavenumber $1.6$, crudely mimicking energy injected by baroclinic instability. The magnitude is determined so that the QG eddy kinetic energy of the vorticity field is one-half:
@@
The fixed-point algorithm described in PolvaniEtAl_94 is then performed to solve for the well-balanced divergence and height fields to match the vorticity fields.
Vorticity asymmetry emerges from the evolution of the shallow water model. We evolve the simulation pseudospectrally using Dedalus (BurnsEtAl_20) for 500 nondimensional time. The horizontal resolution is $1024\times 1024$ Fourier modes. We dissipate the small-scale for numerical stability using a fourth-order hyper-dissipation just large enough to absorb small-scale noise, as diagnosed by the vorticity spectrum
$\nu\nabla^{4}(\cdot)\quad\mbox{with {non-dimensional value}}\quad\nu=5.5\times 10^{-6}.$ (32)
on $u,v,h$ for the shallow water simulation and on $q$ for the SWQG+1 simulation. Time-stepping uses the 3rd-order 4-stage diagonally-implicit+explicit Runge-Kutta (DIRK+ERK) scheme (AscherEtAl_97). The timestep size is determined to have a CFL number of 0.5.
@@ 2.2.2 Simulation results
Ensemble statistics are appropriate for studying the emergent vorticity asymmetry of the shallow water model and whether the SWQG+1 model can capture it. The left of Figure 2 shows the time evolution of vorticity skewness of the size ten ensemble of simulations of the shallow water model and SWQG+1, for $\varepsilon=0.1$. The initial conditions are constructed to have symmetric vorticity, and indeed, the vorticity skewness starts at zero for all simulations. All simulations’ vorticity skewness trends towards negative. The spread between members of the ensembles becomes larger over time, as expected for the chaotic dynamics of shallow water turbulence. However, the ensemble means of vorticity skewness of the two models agree remarkably well during the entire time series. They lie within $1/\sqrt{10}$ times the ensemble standard deviation from each other for most of the time series. We use this range since it is the standard scaling of error of a Monte-Carlo estimate of the ensemble mean. The close agreement indicates strongly that SWQG+1 model can capture the mechanism that leads to vorticity asymmetry in the shallow water model (AraiYamagata_94; GravesEtAl_06).
@@
The agreement extends to all Rossby numbers explored. The right of Figure 2 shows the ensemble mean and $1/\sqrt{10}$ times the ensemble standard deviation of vorticity skewness for the shallow water model and SWQG+1 at nondimensional time $t/T=200$. We see that vorticity skewness becomes more negative as the Rossby number gets larger. PolvaniEtAl_94 shows a similar trend except with the Froude number, which is equal to the Rossby number for all our simulations where the Burger number is one.
@@
which is equal to the Froude number squared. This is consistent with the fact that vorticity skewness is proportional to the Froude number (PolvaniEtAl_94).
@@
The freely evolving simulation in the shallow water model is a simple setting where we can investigate the properties of the SWQG+1 model. We have not been able to show that the SWQG+1 model conserves energy. Instead, we diagnose the energy and show that it mimics the behavior of the energy of the shallow water model. The left of Figure 4 shows the time series of the total energy as well as the EKE and APE components (9) for the evolution of the $\varepsilon=0.1$ ensemble. The EKE decreases as it is converted to the APE, a signature of the inverse cascade. The total energy is well-conserved such that more than 90% of it is in the final state. The SWQG+1 total energy evolution closely follows the one for the full shallow water model, and this is true for all Rossby numbers explored (not shown). However, it is noticeable that the SWQG+1 loses more energy, compared to the shallow water model. The shallow water model dissipates the physical variables separately, while the SWQG+1 model dissipates the PV. Dissipation in the shallow water model does not translate to dissipation of the PV, because of the nonlinearity in the shallow water PV (6). This is well-known for the Ertel PV of the Boussinesq system, where the dissipation of the physical variables translates to a source term in the PV equations (HaynesMcIntyre_87). We will not explore the possibility of including source terms in the PV equation in this work any further. The total energy is minimally dissipated, and the dissipation effects are not significant.
@@ 3.1 The two-layer shallow water system and the derivation of the two-layer SWQG+1 model
$\qquad+\{\varepsilon\}^{2}\left[\left\{\textit{Bu}^{-1}\right\}^{2}\frac{fh_{i}^{2}}{H^{2}}-\left\{\textit{Bu}^{-1}\right\}\frac{(\zeta_{i}+\left\{\gamma\right\}\beta y)h_{i}}{H}\right]+O\left(\{\varepsilon\}^{3}\right)$ (40c)
@@
$\mathcal{L}\begin{pmatrix}\Phi_{1}^{0}\vskip 8.53581pt\\ \Phi_{2}^{0}\end{pmatrix}=\begin{pmatrix}[l]\nabla^{2}\Phi_{1}^{0}+\left\{\textit{Bu}^{-1}\right\}\frac{f^{2}}{g^{\prime}H}(\Phi^{0}_{2}-\Phi^{0}_{1})-\left\{\textit{Bu}^{-1}\right\}\frac{f^{2}}{gH}\Phi_{1}^{0}\vskip 2.84526pt\\ \nabla^{2}\Phi_{2}^{0}+\left\{\textit{Bu}^{-1}\right\}\frac{f^{2}}{g^{\prime}H}(\Phi^{0}_{1}-\Phi^{0}_{2})\end{pmatrix}=\begin{pmatrix}[l]q_{1}\vskip 8.53581pt\\ q_{2}\end{pmatrix}-\left\{\gamma\right\}\beta y.$ (42)
@@
$\mathcal{L}\begin{pmatrix}\Phi_{1}^{1}\vskip 8.53581pt\\ \Phi_{2}^{1}\end{pmatrix}=\begin{pmatrix}[l]C_{q,1}-\left[\left\{\textit{Bu}^{-1}\right\}^{2}\frac{f}{H^{2}}\left(\frac{f}{g^{\prime}}(\Phi^{0}_{1}-\Phi^{0}_{2})+\frac{f}{g}\Phi^{0}_{1}\right)^{2}\right.\\ \left.\qquad-\left\{\textit{Bu}^{-1}\right\}\frac{1}{H}(\nabla^{2}\Phi^{0}_{1}+\left\{\gamma\right\}\beta y)\left(\frac{f}{g^{\prime}}(\Phi^{0}_{1}-\Phi^{0}_{2})+\frac{f}{g}\Phi^{0}_{1}\right)\right]\vskip 2.84526pt\\ C_{q,2}-\left[\left\{\textit{Bu}^{-1}\right\}^{2}\frac{f}{H^{2}}\left(\frac{f}{g^{\prime}}(\Phi^{0}_{2}-\Phi^{0}_{1})\right)^{2}\right.\\ \left.\qquad-\left\{\textit{Bu}^{-1}\right\}\frac{1}{H}(\nabla^{2}\Phi^{0}_{2}+\left\{\gamma\right\}\beta y)\left(\frac{f}{g^{\prime}}(\Phi^{0}_{2}-\Phi^{0}_{1})\right)\right]\end{pmatrix}$ (43)
@@
$\mathcal{L}\begin{pmatrix}F^{1}_{1}\vskip 8.53581pt\\ F^{1}_{2}\end{pmatrix}$ $=\left\{\textit{Bu}^{-1}\right\}\frac{f}{H}\begin{pmatrix}[l]\left.\left[J(\Phi^{0}_{1,x}+\Phi^{0}_{2,x},\Phi^{0}_{1}-\Phi^{0}_{2})+\left\{\gamma\right\}\beta y(\Phi_{1}^{0}-\Phi_{2}^{0})_{,y}\right]\right/g^{\prime}\\ +\left[J(\Phi^{0}_{1,x},\Phi^{0}_{1})+\left\{\gamma\right\}\beta y\Phi_{1,y}^{0}\right]/g\vskip 5.69054pt\\ \left.\left[J(\Phi_{1,x}^{0}+\Phi_{2,x}^{0},\Phi^{0}_{2}-\Phi^{0}_{1})+\left\{\gamma\right\}\beta y(\Phi_{2}^{0}-\Phi_{1}^{0})_{,y}\right]\right/g^{\prime}\end{pmatrix}$, (44)
@@
$\mathcal{L}\begin{pmatrix}G^{1}_{1}\vskip 8.53581pt\\ G^{1}_{2}\end{pmatrix}$ $=\left\{\textit{Bu}^{-1}\right\}\frac{f}{H}\begin{pmatrix}[l]\left[J(\Phi^{0}_{1,y}+\Phi^{0}_{2,y},\Phi^{0}_{1}-\Phi^{0}_{2})+\left\{\gamma\right\}\beta y(\Phi_{1}^{0}-\Phi_{2}^{0})_{,x}\right]/g^{\prime}\\ +\left[J(\Phi^{0}_{1,y},\Phi^{0}_{1})-\left\{\gamma\right\}\beta y\Phi_{1,x}^{0}\right]/g\vskip 5.69054pt\\ \left[J(\Phi_{1,y}^{0}+\Phi_{2,y}^{0},\Phi^{0}_{2}-\Phi^{0}_{1})+\left\{\gamma\right\}\beta y(\Phi^{0}_{2}-\Phi^{0}_{1})_{,x}\right]/g^{\prime}\end{pmatrix}$. (45)
@@ 3.2.1 Set-ups of the simulations
We set our simulation in the atmospheric context, inspired by the work of LambaertsEtAl_12. The idealized two layers atmosphere has $f=10^{4}\text{ s}^{-1}$, $g^{\prime}=2\text{ m/s}^{2}$, $H=10$ km mean thickness, and typical flow speed of $U=5\text{ m/s}$. If we nondimensionalize with a length scale of $L=500$ km, the nondimensional number will be $Bu=8$ and $\varepsilon=0.1$. The domain is doubly-periodic with $Ly=32\times L$ and $Lx=64\times L$.
@@ 3.2.2 Simulation results
The SWQG+1 captures the vorticity evolution of the baroclinic life cycle to remarkable accuracy. Figure 7 displays snapshots of the barotropic ($(\zeta_{1}+\zeta_{2})/2$) and baroclinic ($(\zeta_{1}-\zeta_{2})/2$) vorticity (in color) and divergence (in contour) during the growth of the baroclinic instability. Note that the time of the snapshot is offset. This is because the growth of the instability from small-scale noise happens at different times in the model, due to the chaotic nature of the instability. However, the growth rate is remarkably similar (see Figure 8). The match is indeed impressive. Not only does the SWQG+1 model capture the asymmetric evolution of the cyclonic and anticyclonic part of the baroclinic wave, it is also able to recover the associated divergence fields. Note that one could infer the same divergence, using the $\omega$-equation. The difference of SWQG+1 is that the full velocity participates in the advection, while in SWQG, the higher-order divergence is implicit and is not used in the advection of PV. The divergence is of crucial atmospheric interest as a convergent region in the lower layer corresponds to precipitation. For our example, baroclinic convergence correlates with baroclinic cyclonic vorticity, and baroclinic divergence is large near regions of barotropic strain. On the other hand, barotropic divergence is uniformly small. This is reminiscent of the strain-induced fronts configuration studied in HoskinsBretherton_72. We will explore further strain-induced fronts in Section 3.3. The similarity of the evolution for the two models go beyond this early snapshot. In the supplementary materials, we provide videos of the $\{\varepsilon\}=0.1$ simulations.
@@
Figure 7: The vorticity field of the evolution of the jets for the shallow water model (left) and SWQG+1 model (right). The contour is divergence $\{\varepsilon\}\delta/f$ at [0.1,0.005,0.005,0.1][-0.1,-0.005,0.005,0.1], with the solid being positive. The shallow water snapshot is taken at $t/T=30$ while the SWQG+1 snapshot is taken at $t/T=24$.
@@
The match of the vorticity snapshots translates to the ability of the SWQG+1 to capture the vorticity skewness of the simulation. Figure 8 shows the agreement of the overall trend over a large time ($t/T=100$), for both layers. The vorticity is now skewed towards positive, in comparison to the freely decaying result. This change in vorticity skewness is due to the new effects of the baroclinic mode now included in the two-layer model. The vortex stretching due to the correlation of baroclinic vorticity and divergence drives the initial growth of positive vorticity skewness shown in Figure 8. This effect is entirely missing in the one-layer simulation. The vorticity skewness is an ageostrophic phenomenon. We diagnose the first local maxima of the vorticity skewness, as it is an indicator of the strength of the ageostrophic frontogenesis. Figure 9 shows clearly that this measurement of vorticity skewness is linearly proportional to the Rossby number, and SWQG+1 is effective over a large parameter space.
@@
Our results agree with those of LambaertsEtAl_12 in the upper layer for the onset of instability. However, their lower layer does not have a mean jet while ours does, thus the results differ.
@@ 3.3 Baroclinic strain-driven fronts modeled by SWQG+1
The two-layer simulation reveals that strain-driven fronts are important in the dynamics and contribute to the positive vorticity skewness, which is opposite to the results of our one-layer simulations. This is a new feature of the two-layer shallow water simulations and worth further investigation. Here we study a minimal model of strain-driven fronts by adapting the set-up of HoskinsBretherton_72 to the two-layer model.
@@
$\mathcal{L}\begin{pmatrix}G^{1}_{1}\vskip 8.53581pt\\ G^{1}_{2}\end{pmatrix}$ $=\left\{\textit{Bu}^{-1}\right\}\frac{2\alpha f}{g^{\prime}H}\begin{pmatrix}[l](\Phi^{0}_{2}-\Phi^{0}_{1})_{,y}\vskip 5.69054pt\\ (\Phi^{0}_{1}-\Phi^{0}_{2})_{,y}\end{pmatrix}$. (50)
@@
The HoskinsBretherton_72 model is posed in an infinite domain in $y$, where perturbation velocities decay at infinity. This implies boundary conditions for the potentials:
@@
$\delta_{1}=v_{1,y}$ $=-\{\varepsilon\}G_{1,y}\quad\mbox{and}\quad\delta_{2}=v_{2,y}=-\{\varepsilon\}G_{2,y}.$ (53)
overlaid on the height perturbations. In the top layer, a cold filament has cyclonic vorticity, which is correlated with convergence. This correlation is a key feature of fronts and explains the cyclonic bias of vorticity in the two-layer simulations by appealing to vortex stretching: baroclinic cyclonic vorticity is correlated with baroclinic negative divergence. Note that the above equation is equivalent to the $\omega$-equation”. Therefore, it is implicit in the SWQG model. However, SWQG does not use this ageostrophic velocity in its advective evolution. It does not capture the vorticity skewness evolution.
Finally, a comparison with the surface buoyancy gradient driven Boussinesq frontogenesis of HoskinsBretherton_72 is warranted. In the layered shallow water set-up, the convergence will smooth away the height perturbation and release the stored APE. Shallow water fronts would be less strong with no finite time singularity, compared to the Boussinesq case.
Figure 10: Divergence of a cold filament driven by a barotropic strain, modeled by a two-layer rigid lid SWQG+1 model. The domain is divided vertically into two layers by the height perturbation. The color shows divergence values. This is a zoomed-in view showing y/L=[3,3]y/L=[-3,3]